BOB'S FATHER IS 3 TIMES OLD AS BOB. 4 YEARS AGO HE WAS 4 TIMES OLDER. HOW OLD IS BOB?
Answer by uksraj(9)

let bob's father current age is x;and that of bob is y
age at four years ago i.e of father=x-4 & of bob =y-4
given 3y=x I
(x-4)=4*(y-4) II
putting value of equ I in II
we have 3y-4=4y-16 i.e y=12 & x=3x12=36